Without calculating the determinant, inspect the following:


(7,1) 1 0 0 0

0 1 0 0

0 0 1 0

0 0 0 −2


(7.2) 1 0 0 0

0 1 0 0

0 0 0 1

0 0 1/4 0


1
Expert's answer
2021-06-03T16:50:12-0400

(7.1)

"\begin{vmatrix}
 1&0&0&0\\0&1&0&0\\0&0&1&0\\0&0&0&-2
\end{vmatrix}"


Since it is a diagonal matrix.

So, the determinant can be calculated by inspection(by multiplying the diagonal elements)


"Determinant = 1\times 1\times 1\times (-2)\\\boxed{Determinant=-2}"



(7.2)

"\begin{vmatrix}
 1&0&0&0\\0&1&0&0\\0&0&0&1\\0&0&1/4&0
\end{vmatrix}"


Interchanging third and fourth row , which is a valid transformation with respect to the determinant (it will leave it unchanged), you will get:


"\begin{vmatrix}
1&0&0&0\\0&1&0&0\\0&0&-1/4&0\\0&0&0&1
\end{vmatrix}"



Since it is a diagonal matrix.

So, the determinant can be calculated by inspection(by multiplying the diagonal elements)


"Determinant = 1\times 1\times( -1/4)\times 1\\\boxed{Determinant=-\dfrac{1}{4}}"